OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (4, –6); m = 3/5 a. y - 6 = (3/5)(x + 4) b. y + 6 = (3/5)x - 4 c. y - 4 = (3/5)(x + 6) d. y + 6 = (3/5)(x - 4)

OpenStudy (dumbcow):

point slope formula is: (y-y1) = m(x-x1) m = slope point has x,y coordinates of (x1,y1) substitute values into point slope formula

OpenStudy (dumbcow):

it will be exact same for all the problems you posted

OpenStudy (dumbcow):

so what are you getting for an answer